### v1

@@ 3 Discussion
This study has several limitations. First, we evaluated feasibility solely from clinicians’ perspectives using a standardized patient, without input from real patients. Second, although we hypothesized that enhanced personalization would improve treatment relevance and effectiveness—consistent with prior literature [golas2021predictive, hwang2023personalization, yardley2015person]—this requires confirmation in longitudinal prospective trials. Third, the study focused on a single patient profile and a small therapist cohort; broader inclusion of diverse conditions and a larger clinician sample will be needed to assess whether LLMs can faithfully translate heterogeneous treatment strategies into software with appropriate monitoring logic. Finally, in the absence of benchmark data, we conducted a post hoc comparison against a simulated external benchmark (a generalized software template). Demonstrating superiority in personalization will require a prospective, two-arm trial directly comparing LLM-generated interventions with existing digital rehabilitation programs.
In summary, we introduce a digital prescription paradigm in which clinicians use an LLM to generate intervention software after clinical encounters, informed by individual patients’ needs. This contrasts with the prevailing approach, in which clinicians retrofit pre-programmed software libraries authored prior to clinical encounters. To our knowledge, the prospective feasibility of clinician-directed, LLM-based software generation has not been evaluated previously in any healthcare domain. Our results demonstrate feasibility and therapist acceptability in a standardized-patient setting and suggest applicability beyond rehabilitation. These findings merit evaluation in larger, more diverse clinical settings with real patients and longer-term outcomes.
@@ 4.1 Participants
We recruited 20 physical and occupational therapists through emails to rehabilitation centers or clinics within the United States. Inclusion criteria were that participants be registered and licensed physical or occupational therapists in the United States and have practiced for at least two years. Written informed consent was obtained before enrollment. This study was reviewed and approved by the institutional review board (UC Berkeley IRB protocol ID: 2025-01-18162). The study lasted for 1 hour per therapist. Each therapist received USD \$50 in compensation.
@@ 4.3 LLM Software Generation
As shown in Fig. 1, we provided the following inputs to the LLM in addition to the therapist’s exercise descriptions: (i) documentation on syntax and semantics of the programming language, (ii) an API library, (iii) a coding guideline for software generation, and (iv) example programs pairing sample exercise prescriptions with corresponding exemplary software.
@@ 4.5 Experiment Setup
Exercise goals. We prepared 10 upper extremity exercise goals, each with a corresponding worksheet adapting from Shirley Ryan Ability Lab’s upper extremities exercises (e.g., [shirley_ryan]), in consultation with licensed therapists and neurologists on the study team (refer to Appendix B). Each therapist was randomly assigned 2 goals, resulting in 4 therapists per goal (refer to Appendix D). This study design ensured that all 10 goals were covered and allowed us to examine potential variations in personalization of exercises across therapists.
@@ 4.6.1 Pre-labeled Evaluation for Monitoring Accuracy
To evaluate the intervention software’s monitoring capability, therapists conducted a pre-labeled assessment. Before observing the intervention software interacting with the standardized patient, each therapist was presented with a table (Fig. 4, left) listing the sequence of instructions that the software would deliver. To rigorously assess monitoring accuracy, therapists pre-labeled which instructions the patient should complete and which should remain incomplete. This approach ensured that the evaluation reflected the software’s actual monitoring capability and prevented artificially inflated results. For example, if the patient happened to complete all steps, the software could achieve 100% accuracy by marking every instruction as complete without truly monitoring. Pre-labeling also created a balanced and known distribution of outcomes, enabling the calculation of accuracy, sensitivity, and specificity. The proportion of instructions that therapists pre-labeled as incomplete was 36.3%.
4.6.2 Comparison of Existing vs Proposed DHI Paradigms
In the existing digital intervention paradigm for remotely instructing and monitoring patients, physical and occupational therapists typically select exercise treatments from libraries of parametrized intervention software that are programmed prior to clinical encounters. These software modules encode exercise instructions derived from standard exercise worksheets, while exposing a limited set of parameters (for example, number of repetitions) that clinicians can adjust for individual patients.
@@ 4.6.2 Comparison of Existing vs Proposed DHI Paradigms
In our study, we constructed a proxy for this status quo paradigm by preparing ten exercise worksheets corresponding to ten common exercise goals (Sec. 4.5). The instructions contained in these ten worksheets were taken to represent the fixed instruction sets used in intervention software. In this paradigm, the textual instructions remain fixed, and therapists can only adjust predefined parameters. Because there is no standardized schema for how existing DHIs parametrize exercises, we derived a generalized parameter set by aggregating the parameters used in digital rehabilitation trials in physical and occupational therapy over the past ten years. This yielded a generalized DHI template that approximates the current status quo. The template parameters included exercise side (left/right/bilateral) [shebib2019lowback, pak2023shoulderpain], number of repetitions [cui2023lowback, malliaras2020intel, mecklenburg2018kneepain], pose hold time [weise2022backpain], and predefined difficulty or goal options (e.g., ranges of motion or target locations) [prieto2024activehip, pak2023shoulderpain, gauthier2022vigorous, bailey2020chronicpain, biebl2021app].
@@
After each clinical encounter with the standardized patient, we asked therapists to prescribe personalized exercises for their assigned exercise goals (Appendix B,D); each exercise goal was assigned to four different therapists to observe variations in personalized exercises (Sec. 4.5). We then examined, for each of the 40 personalized prescriptions, whether its exercise instructions could be represented within the existing paradigm by (i) selecting one of the ten worksheet-based instruction sets and (ii) adjusting only the parameters allowed by the generalized DHI template. If a therapist’s prescription could not be expressed under these constraints, we considered it not translatable into intervention software under the existing template-based paradigm.

### v2

@@ 3 Discussion
This study has several limitations. First, we evaluated feasibility solely from clinicians’ perspectives using a standardized patient, without input from real patients. Second, although we hypothesized that enhanced personalization would improve treatment relevance and effectiveness—consistent with prior literature [golas2021predictive, hwang2023personalization, yardley2015person]—this requires confirmation in longitudinal prospective trials. Third, the study focused on a single patient profile and a small therapist cohort; broader inclusion of diverse conditions and a larger clinician sample will be needed to assess whether LLMs can correctly translate heterogeneous treatment strategies into software with appropriate monitoring logic. Finally, in the absence of benchmark data, we conducted a post hoc comparison against a simulated external benchmark (a generalized software template). Demonstrating superiority in personalization will require a prospective, two-arm trial directly comparing LLM-generated interventions with existing DHI platforms.
In summary, we introduce a novel digital prescription paradigm in which clinicians use an LLM as a constrained translator to generate intervention software after clinical encounters, once individual patients’ needs are fully understood. This reverses the prevailing approach, in which clinicians must retrofit prescriptions into pre-programmed software libraries authored before encounters. In a prospective feasibility study, this paradigm enabled near-complete translation of personalized prescriptions into software, high accuracy of instruction delivery, and acceptable preliminary monitoring performance, with most therapists judging the system safe and expressing willingness to adopt it in practice. To our knowledge, this is the first prospective evaluation of clinician-directed, LLM-based software generation in healthcare. These findings position LLMs as a practical tool for scaling personalized DHIs and motivate longitudinal prospective trials with real-patient populations.
@@ 4.1 Participants
We recruited 20 physical and occupational therapists through emails to rehabilitation centers or clinics within the United States. Inclusion criteria were that participants be registered and licensed physical or occupational therapists in the United States and have practiced for at least two years. Written informed consent was obtained before enrollment. This study was reviewed and approved by the institutional review board (UC Berkeley IRB protocol ID: 2025-01-18162). The study lasted for 1 hour per therapist. Each therapist received USD 50 in compensation.
@@ 4.3 LLM Software Generation
As shown in Fig. 1, we provided the following system prompt to the LLM in addition to the therapist’s exercise descriptions: (i) documentation on syntax and semantics of the programming language, (ii) an API library, (iii) a coding guideline for software generation, and (iv) example programs pairing sample exercise prescriptions with corresponding exemplary software. The system prompt used for the study can be found in Appendix A.
@@ 4.5 Experiment Setup
Exercise goals. We prepared 10 upper extremity exercise goals, each with a corresponding worksheet adapting from Shirley Ryan Ability Lab’s upper extremities exercises (e.g., [shirley_ryan]), in consultation with licensed therapists and neurologists on the study team (refer to Appendix B). Each therapist was randomly assigned 2 goals, resulting in 4 therapists per goal (refer to Appendix D). This study design ensured that all 10 goals were covered and allowed us to examine potential variations in personalization of exercises across therapists. In total, 40 personalized exercise prescriptions are created by therapists (refer to Appendix C).
@@ 4.6.1 Pre-labeled Evaluation for Monitoring Accuracy
To evaluate the software’s monitoring capability, therapists completed a pre-labeled assessment. Before observing the software interact with the standardized patient, each therapist viewed a table listing the sequence of instructions the software would deliver (Fig. 4, left). For each instruction, therapists indicated in advance whether the patient should complete it or leave it incomplete. This design ensured that monitoring accuracy reflected the software’s detection performance rather than the patient’s actual behavior. For example, if the patient completed every step, a system that arbitrarily marked all instructions as complete would appear to be 100% accurate without actually monitoring. Pre-labeling created a known mix of complete and incomplete instructions, allowing us to compute accuracy, sensitivity, and specificity; overall, 36.3% of instructions were pre-labeled as incomplete.
4.6.2 Comparison of Existing and Proposed DHI Paradigms
Current commercial DHIs for remote exercise instruction typically provide libraries of pre-programmed intervention software that instantiate standard exercise worksheets targeting broad patient populations. These modules encode fixed textual instructions (for example, generic upper extremity strengthening programs) and expose a limited set of parameters—such as number of repetitions, difficulty level, or target side—to tailor treatment to individual patients. In routine use, clinicians choose a module that roughly matches a patient’s impairment and then adjust these parameters. However, they generally cannot modify the underlying instructions to incorporate patient-specific contingencies (for example, alternative steps when a patient cannot perform a motion) or contextual details (for example, household objects, home layout, or work schedule). Because existing systems are heterogeneous and proprietary, there is no standardized benchmark or parameter schema describing how such exercise libraries are structured.
@@ 4.6.2 Comparison of Existing and Proposed DHI Paradigms
To enable a systematic comparison with this prevailing digital intervention paradigm, we constructed a generalized DHI template based on standardized exercise worksheets and recent digital rehabilitation trials in physical and occupational therapy. We first derived a generalized set of exercise parameters by aggregating those reported in digital rehabilitation trials over the past ten years, including exercise side (left/right/bilateral) [shebib2019lowback, pak2023shoulderpain], number of repetitions [cui2023lowback, malliaras2020intel, mecklenburg2018kneepain], pose hold time [weise2022backpain], and predefined difficulty or goal options (for example, ranges of motion or target locations) [prieto2024activehip, pak2023shoulderpain, gauthier2022vigorous, bailey2020chronicpain, biebl2021app]. We then instantiated this generalized template using ten exercise worksheets adapted from existing standardized worksheets, each corresponding to a common exercise goal (Sec. 4.5). In this generalized DHI template, the worksheet instructions represent the fixed exercise content, and clinicians are allowed to adjust only the template parameters, mirroring how current DHIs permit parameter tuning without altering the underlying instructions.
@@
After each clinical encounter with the standardized patient, therapists prescribed personalized exercises for their assigned exercise goals (Appendix B,D); each goal was assigned to four different therapists to capture variation in personalized design (Sec. 4.5). Our intent was to quantify how these post-encounter, therapist-authored prescriptions deviate from the standard worksheet–based instructions that underpin current DHIs, and to assess whether those deviations could be accommodated by parameter adjustments alone. For each of the 40 personalized prescriptions, we therefore examined whether it could be retrofitted into the generalized DHI template for the corresponding exercise goal by (i) selecting the associated worksheet-based instruction set and (ii) adjusting only the allowed parameters. If a prescription required changing the textual instructions (for example, adding contingencies or patient-specific context) rather than just tuning parameters, we considered it not translatable into intervention software within the current digital intervention paradigm.
@@ 4.7 Data Availability & Reproducibility
5 Funding
This work was supported by the UC Noyce Initiative’s Computational Health Program, which funds innovative research at the intersection of computing, digital transformation, medicine, and the health sciences, and supported the collaboration between UC Berkeley and UC San Francisco. Authors affiliated with Stanford University contributed to this work without dedicated funding support.
6 Conflict of Interest
All of the authors declared that they have no conflicts of interest.
7 Author Contributions
Edward Kim (E.K.) and Yuri Cho (Y.C.) co-led the research. E.K. identified the core problem, co-designed and conducted the study, led the overall technical direction, architected and implemented the LLM-based intervention software system, coordinated system development, and wrote the Results and Methods sections. Y.C. led the clinical study design, formulated the study hypotheses, conducted the study, and drafted the Introduction and Discussion sections. José Lima (J.L.) refined the study design and performed the statistical analyses. Julie Muccini (J.M.) refined the study design, particularly the definition of exercise goals and worksheets, and iteratively tested the system, providing feedback from an occupational therapy perspective. Jenelle Jindal (J.J.) contributed to refining the study design. Allison Scheid (A.Sc.) tested the LLM-based tool and provided feedback from a physical therapy perspective. Erik Nelson (E.N.) designed and implemented the back-end server used in the study. Sun Min Kim (S.K.) implemented the database and cloud-based program synthesis routines. Seong Hyun Park (S.P.) led the design and implementation of the augmented reality (AR)–based patient-facing application and its integration with the tele-rehabilitation platform, enabling remote therapist–patient sessions. Alton Sturgis (A.St.) implemented the video call functionality of the tele-rehabilitation platform in the AR headset. Caesar Li (C.L.) implemented real-time body pose estimation to detect and analyze movements. Jackie Dai (J.D.) developed algorithms to capture images of body movements via AR headset cameras for real-time analysis. Yash Prakash (Y.P.) designed and implemented the multiplayer Unity server that enabled clinician–patient interaction within the tele-rehabilitation platform. Daniel He (D.H.) integrated Scenic with Unity to support instruction and monitoring of patient body movements. Wiktor Rajca (V.R.) developed algorithms to parse movement trajectories into interpretable joint angles for real-time analysis. Yuchen Zeng (Y.Z.) led development of the web portal used by clinicians to conduct sessions and prescribe exercises. Liwen Sun (L.S.) designed and implemented the evaluation schema therapists used to assess patients via the portal. Isabella Hu (I.H.) architected the portal’s video call functionality and workflow for recording and parsing sessions to generate exercise worksheets using LLMs. Hongxuan Wu (H.W.) led the visual design of the portal. Dr. Cathra Halabi (C.H.) refined the study design and provided clinical oversight from a neurology perspective. Dr. Maarten Lansberg (M.L.) refined the study design, contributed to interpretation of results, and advised on statistical analysis. Professors Sanjit Seshia (S.S.) and Bjoern Hartmann (B.H.) advised the project from the computer science side, helped refine the study design, supervised system development, and served as technical co–principal investigators. Dr. Halabi and Dr. Lansberg served as clinical principal investigators. All authors reviewed, edited, and approved the final manuscript.